OpenStudy (anonymous):

Which of the following sets of numbers is a Pythagorean triple? 14, 40, 50 7, 24, 31 7, 20, 27 14, 48, 50

OpenStudy (anonymous):

Do yah know how to find a pythagorean triple? ;o

OpenStudy (anonymous):

Im having a brain fart so, no lol

OpenStudy (anonymous):

use a^2+b^2=c^2 where c is the largest number in the set. Then just plug in until you get something that is true. I'll make an example for first choice below.

OpenStudy (anonymous):

WELL. You do the \[a^2+b^2=c^2\] and then you have to input the numbers, to find out which actually comes out to equal c^2. id that made sense

OpenStudy (anonymous):

A and B are the lowest numbers and C is the biggest number. Just use \[A ^{2}+B ^{2}=C ^{2}\] Plug the numbers in. The one that equals is your answer.

OpenStudy (anonymous):

14, 40, 50 here the 50 is largest number and so you plug it in c. the other 2 dont matter the order because of the property of addition. 14^2+40^2=50^2 you evaluate and if true that is your triple. Do it for each choice until you find one that is true.

OpenStudy (anonymous):

Okay I think I got it.

OpenStudy (anonymous):

The answer is D. Or the last one. In case you want to check.

OpenStudy (anonymous):

Thank you!